OpenStudy (anonymous):

Suppose that the demand and price for strawberries are related by p equals D of q equals 5 minus 0.25 q where p is the price (in dollars) and q is the demand (in hundreds of quarts). Also suppose the price and supply of strawberries are related by p equals S of q equals 0.25 q Find the equilibrium quantity.

OpenStudy (anonymous):

I have not seen a problem like this in years, but this may be as simple as stating D(q) = S(q); but I am easily incorrect

OpenStudy (anonymous):

Here's what we know. D(q)=5-0.25q S(q) = 0.25q You have demand meeting supply when D(q)=S(q) that is to say 5-0.25q = 0.25q Solve for q. Of course, I too am easily incorrect! No, really, no kidding. I have no idea.